OpenStudy (anonymous):

Which of the following is the third term in the expansion of (8 – x)^3?

OpenStudy (anonymous):

3 * 8 * x^2 = 24x^2

OpenStudy (anonymous):

Using the identity.

OpenStudy (anonymous):

(8-x)^3 (8-x)(8-x)(8-x) (64-16x+x^2)(8-x) (512-128x+8x^2-64x+16x^2-x^3) -x^3+24x^2-192x+512 So the third term is -192x

OpenStudy (anonymous):

You can use the identity. According to the identity, the third term is 3xy^2. So the third term is 24x^2.

OpenStudy (anonymous):

Yellowranger has given us the second term instead of the third one.......

OpenStudy (anonymous):

Use the binomial theorem: \[(a+b)^n=\sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k} b^{k}\] Replace n with 3 giving: \[(a+b)^3=\sum_{k=0}^{3}\left(\begin{matrix}3 \\ k\end{matrix}\right)a^{3-k} b^{k}=a^3+3a^2b+3ab^2+b^3\]

OpenStudy (anonymous):

So am I right?

OpenStudy (anonymous):

Once again pratu has been cooler than me at answering math questions. I counted to three from left to right like the misinformed miscreant that I am . Thank you, pratu.